OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, -1).

OpenStudy (anonymous):

reqd line is x-3y=k where k is constant now as it passes through (6,-1) thus putting these in the obtained eq. we have 6+3=k or k=9 thus reqd. line is x-3y=9

OpenStudy (anonymous):

thankyou sooo much